OpenStudy (anonymous):

An investor invested a total of $2,300 in two mutual funds. One fund earned a 6% profit while the other earned a 3% profit. If the investor's total profit was $123.00, how much was invested in each mutual fund?

OpenStudy (anonymous):

u would need to let x = the amt invested at the 6% amt.. and let 2300-x = the amount invest at the 3%

OpenStudy (anonymous):

distribute and solve for x.

OpenStudy (anonymous):

.06x + .03x(2300 -x) = 123

OpenStudy (anonymous):

so x = 1800. then go back in and sub.. 2300 - 1800 = 500. so he invested 1800 @6% and 500 @ 3%